# Discovery of a dwarf planet candidate in an extremely wide orbit: 2017 OF201

### v1

@@ 4.1 Possible origin of 2017 OF201
where $a_{\rm N}$ is the semi-major axis of Neptune and $q$ is the perihelion distance of 2017 OF201. This implies that 2017 OF201 cannot reach its current large semi-major axis through perturbations from Neptune alone given the age of the solar system. Instead, its high-$q$, large-$a$ orbit likely reflects the influence of external torques from stellar encounters or the Galactic tide, which act during or after an initial scattering event(Duncan et al., 1987; Bannister et al., 2017; Gladman & Volk, 2021; Batygin et al., 2021).
@@
2017 OF201 may have first been scattered outward by Neptune while retaining a low perihelion. Once at large $a$, the Galactic tide becomes dynamically important, gradually increasing $q$ and detaching the orbit from Neptune. As the perihelion rises into a region where planetary perturbations are weak, the object undergoes slow angular momentum oscillation and reaches a configuration like that observed today (see Figure 6 for a schematic illustration).
To qualitatively test the plausibility of this migration scenario, we perform backward N-body integrations over a 3.5 Gyr timescale. We simulate 100 realizations of the system, each initializes 2017 OF201 with orbital parameters drawn from a multivariate normal distribution based on the best-fit values and uncertainties in Table 1.
Figure 3 shows a representative backward evolution of 2017 OF201’s orbit, illustrating the final stage of the migration pathway discussed above. Once 2017 OF201 resides on a high-$q$, large-$a$ orbit, the Galactic tide gradually reduces its angular momentum, $J\propto\sqrt{a(1-e^{2})}$, leading to a slow decrease in $q$ (Heisler & Tremaine, 1986). Once the perihelion enters the planetary region, weak perturbations – primarily from Neptune – scatter 2017 OF201 into a more tightly bound orbit. While not definitive, this outcome supports the dynamical viability of a distant scattering-disk origin followed by perihelion lifting and subsequent inward migration.
Figure 3: A representative backward integration of 2017 OF201’s orbit over 3.5 Gyr, showing the semi-major axis $a$ (top) and perihelion distance $q$ (bottom) as functions of time. The gradual decrease in the orbital angular momentum driven by the Galactic tide leads to a slow lowering of $q$. At later times, 2017 OF201’s perihelion enters the planetary region, where weak interactions scatter it inward.
4.2 Implications for the hypothetical Planet X
As shown in Figure 2, the longitude of perihelion of 2017 OF201, $\varpi=306^{\circ}$, lies outside the clustering region near $\varpi\approx 60^{\circ}$ observed among other extreme TNOs. This distinction raises the question of whether 2017 OF201 is dynamically consistent with the Planet X hypothesis, which suggests that a distant massive planet shepherds TNOs into clustered orbital configurations.
Siraj et al. (2025) computed the most probable orbit for a hypothetical Planet X by requiring that it both reproduces the observed clustering in the orbits of extreme TNOs and preserves the long-term stability of those TNOs that are already stable in the absence of Planet X. To assess how such a planet would affect 2017 OF201, we perform forward N-body integrations including a planet with mass $m_{p}=4.1,M_{\oplus}$, semi-major axis $a_{p}=296$ au, eccentricity $e_{p}=0.339$, inclination $i_{p}=4.27^{\circ}$, and longitude of perihelion $\varpi_{p}=242^{\circ}$ (Siraj et al., 2025). We compare this to a control simulation that excludes Planet X.
@@ 4.2 Implications for the hypothetical Planet X
Figure 4 shows that, in the absence of Planet X, 2017 OF201 remains on a stable orbit for at least 1 Gyr under the influence of the known giant planets and the Galactic tide. In contrast, when Planet X is included, gravitational interactions lead to close encounters with Neptune that eject 2017 OF201 from the solar system in $\sim$0.1 Gyr.
Figure 4: Forward dynamical evolution of 2017 OF201 with (red) and without (blue) the influence of a hypothetical Planet X with orbital parameters from Siraj et al. (2025). Without Planet X, the orbit of 2017 OF201 remains stable for more than 1 Gyr under the influence of the known giant planets and the Galactic tide. In contrast, in the presence of Planet X, 2017 OF201 experiences close encounters with Neptune and is ejected from the solar system in $\sim$0.1 Gyr.
These results suggest that the existence of 2017 OF201 may be difficult to reconcile with this particular instantiation of the Planet X hypothesis. While not definitive, 2017 OF201 provides an additional constraint that complements other challenges to the Planet X scenario, such as observational selection effects and the statistical robustness of the observed clustering.
5 Conclusion
@@
We report the discovery of an exotic trans-Neptunian object 2017 OF201 from the archival data of DECaLS. Its orbit is extremely wide and eccentric ($a=838$ au, $q=44.9$ au, $i=16.2^{\circ}$). It is currently located at a distance of 90.5 au and is among the top ten most distant solar-system objects observed in the optical. Furthermore, it has the second brightest absolute magnitude $H_{V}=3.5$ among known TNOs with wide orbits ($a>$ 80 au). When assuming a typical albedo of 0.15, the diameter of 2017 OF201 is estimated to be 700 km, large enough to qualify as a dwarf planet.
The discovery of 2017 OF201 suggests a population behind it with hundreds of objects possessing similar properties, because the probability for 2017 OF201 to be close enough and detectable is only 0.5% given its wide and eccentric orbit. Estimated from the large size of 2017 OF201, the total mass of that population is potentially 1% of Earth’s mass, which is a significant fraction of the scattering disk.
@@
We show that the orbit of 2017 OF201 is shaped by both Neptune’s scattering and the Galactic tide at Gyr timescale (Figures 3 and 6), thus its orbit represents an overlap between the scattering disk and the inner Oort cloud.
The orbit of 2017 OF201 also presents an intriguing challenge to the Planet X / Planet 9 hypothesis. Previous discoveries of TNOs with extremely wide and elongated orbits suggest an apparent clustering in the longitude of perihelia $\varpi\approx 60^{\circ}$, motivating the idea that an undetected planet may be shepherding the extreme TNOs. However, with $\varpi$ = 306, the orbit of 2017 OF201 is an outlier of that clustering (Figure 2) and likely unstable to gravitational influence from this hypothesized planet. Our $N$-body simulations suggest that the presence of the Planet X / Planet 9 that produces the clustering will cause ejection of 2017 OF201 in a short timescale around 0.1 Gyr (Figure 4). Nevertheless, more investigation is required to decisively rule out that hypothesis.
@@
Figure 5: Cutout images of all 19 detections from both CFHT and DECam. Each image is 20 arcsec on a side. The detection time corresponds to the middle of each exposure. CFHT data are deeper and have a much better seeing than the DECam data. Our detection spans a long time baseline from Aug 2011 to Oct 2018, enabling precise determination of its orbit.
Figure 6: Left: TNOs with wide orbits ($a>$150 au), 2017 OF201 has the second largest brightest among this population. Right: Schematic of a possible migration pathway for 2017 OF201. The object may have been scattered to large semi-major axis ($a$) while maintaining a low perihelion distance ($q$). Galactic tidal torques then slowly increase $q$, detaching the orbit from Neptune.
@@ Appendix A Astrometry and Photometry
We perform astrometric and photometric measurements of the object using both CFHT/MegaCam and DECam imaging data. Cutouts of those images with 2017 OF201 centered are shown in Figure 5.
@@
For the detections in CFHT, we retrieve the single-epoch images from the Canada-France-Hawaii Telescope (CFHT)/MegaCam data archive [^12] https://www.cadc-ccda.hia-iha.nrc-cnrc.gc.ca/. These images are all in the $r$-band with 0.5–0.9 arcsec seeing. The instrument signatures (bias, dark, and flat) have been removed in these images, but they do not yet have astrometric solutions and photometric calibrations. We implement a two-step astrometric calibration process: first obtaining an initial solution using astrometry.net (Lang et al., 2010), followed by fine-tuning against Gaia DR3 reference stars (Gaia Collaboration et al., 2021) using SCAMP (Bertin, 2006). This procedure yields a final astrometric solution with an average positional uncertainty of $0.1\arcsec$. For photometric calibration, we first remove the sky background using SExtractor (Bertin & Arnouts, 1996) with a mesh size of $60\arcsec$. We then construct a point spread function (PSF) model using PSFEx (Bertin, 2011) from non-saturated point sources in each image, and perform PSF-fitting photometry using SExtractor. The photometric zeropoint is thus determined by comparing the instrumental magnitudes of non-saturated stars with their Pan-STARRS1 PSF magnitudes. Our object has an average $r$-band magnitude of $r=22.61$ mag in the CFHT data from 2011 to 2012.
@@
The DECam single-epoch images are downloaded from the Legacy Surveys [^13] https://www.legacysurvey.org/rawdata/, which are reduced using the DECam Community Pipeline [^14] https://noirlab.edu/science/data-services/data-reduction-software/csdc-mso-pipelines/pl206. These images are photometrically calibrated using Pan-STARRS1 (Chambers et al., 2016) and have a preliminary astrometric solution. To ensure consistency with our CFHT analysis, we refine the astrometric calibration using SCAMP against the Gaia DR3 catalog. Similar to the CFHT data above, for each DECam image, we subtract the sky background with a mesh size of $60\arcsec$, generate the PSF model, and perform PSF-fitting photometry. The resulting photometric results have higher uncertainties since the DECam images are shallower and have poorer seeing conditions (0.8–2.0 arcsec) than the CFHT images.

### v2

@@ IV.1 Possible origin
where $a_{\rm N}$ is the semi-major axis of Neptune and $q$ is the perihelion distance of 2017 OF201. This implies that 2017 OF201 cannot reach its current large semi-major axis through perturbations from Neptune alone, given the age of the solar system. Instead, its high-$q$, large-$a$ orbit likely reflects the influence of external torques from stellar encounters or the Galactic tide, which act during or after an initial scattering event (Duncan et al., 1987; Bannister et al., 2017; Gladman and Volk, 2021; Batygin et al., 2021).
@@
2017 OF201 may have first been scattered outward by Neptune while retaining a low perihelion. Once at large $a$, the Galactic tide becomes dynamically important, gradually increasing $q$ and detaching the orbit from Neptune. As the perihelion rises into a region where planetary perturbations are weak, the object undergoes slow angular momentum oscillation and reaches a configuration like that observed today (Figure 3). Or, alternatively, a passing star in the birth cluster of the Sun may provide the torque to raise $q$ (Nesvorný et al., 2023).
IV.2 Long-term stability
We investigate the long-term dynamical stability of 2017 OF201 using direct N-body integrations with the ias15 integrator (Rein and Spiegel, 2015) in REBOUND [^13] https://rebound.readthedocs.io/ (Rein and Liu, 2012). Because scattering of distant objects is primarily driven by Uranus and Neptune (Batygin and Brown, 2016; Gerdes et al., 2017), we explicitly include these planets in our simulations. Jupiter and Saturn are incorporated through a static solar quadrupole moment ($J_{2}$), which efficiently captures their secular influence.  [^14] For validation, we also performed computationally intensive integrations that directly included Jupiter and Saturn. The outcomes were qualitatively consistent with our simplified model. Given 2017 OF201’s large aphelion distance ($\sim 1600$ au), we additionally include the vertical component of the Galactic tide (Levison et al., 2001) using the add_force module in REBOUNDx [^15] https://reboundx.readthedocs.io/ (Tamayo et al., 2020).
We draw 100 realizations of 2017 OF201’s orbits from a multivariate normal distribution based on the best-fit orbital parameters and their uncertainties, and integrate each system for 4 Gyr. At the end of the integrations, 2017 OF201 remains bound in more than 80% of realizations, yet in roughly half of cases its semi-major axis varies by a factor of $\geq 2$. Following the criterion of Batygin and Brown (2021), we therefore classify 2017 OF201 as metastable.
@@ IV.2 Long-term stability
Figure 4: Typical clones of dynamical evolution simulation of 2017 OF201, with and without P9. In the absence of P9, 2017 OF201 remains stable over 4 Gyr under the influence of the known giant planets and the Galactic tide (the x-axis is truncated for clarity). By contrast, when P9 is included, 2017 OF201 undergoes close encounters with Neptune and is ejected from the Solar System within $\sim$0.3 Gyr.
IV.3 Implications for the hypothetical Planet Nine
As shown in Figure 2, the longitude of perihelion of 2017 OF201, $\varpi=306^{\circ}$, lies well outside the clustering region near $\varpi\approx 60^{\circ}$ observed among other extreme TNOs. This discrepancy raises the question of whether 2017 OF201 is dynamically compatible with the Planet Nine (P9) hypothesis, which posits that a distant massive planet shepherds TNOs into aligned orbital configurations.
Brown and Batygin (2021) inferred a candidate P9 with mass $m_{p}=6.2,M_{\oplus}$, semimajor axis $a_{p}=382.4$ au, eccentricity $e_{p}=0.20$, inclination $i_{p}=15.6^{\circ}$, and longitude of perihelion $\varpi_{p}=246.7^{\circ}$. More recently, Siraj et al. (2025) derived an updated configuration that simultaneously reproduces the observed orbital clustering and preserves the stability of already stable TNOs, yielding $m_{p}=4.1,M_{\oplus}$, $a_{p}=296$ au, $e_{p}=0.339$, $i_{p}=4.27^{\circ}$, and $\varpi_{p}=242^{\circ}$.
To evaluate the influence of a putative P9 on 2017 OF201, we performed forward N-body integrations including P9 with each of the parameter sets described above, and compared the results against a control case without P9. For each configuration, we fixed all orbital elements of P9 to their mean values and sampled 100 values of the mean anomaly of P9 uniformly across $0,2\pi]$, integrating the systems for 4 Gyr. With the parameters proposed by Brown and Batygin (2021), 56% of realizations resulted in ejection of 2017 OF201, whereas the updated parameters from Siraj et al. (2025) produced a 64% ejection fraction. The corresponding mean ejection timescales are 0.31 and 0.25 Gyr, respectively. Figure 4 shows typical clones of the simulation result.
These results suggest that the orbit of 2017 OF201 is difficult to reconcile with the specific P9 configurations from Brown and Batygin (2021) and Siraj et al. (2025). Therefore, 2017 OF201 offers an additional challenge to the P9 hypothesis, complementing other challenges such as observational selection effects and questions regarding the statistical significance of the reported clustering (Shankman et al., 2017; Bernardinelli et al., 2020a; Napier et al., 2021).
V Conclusion
@@
We report the discovery of a large, exotic trans-Neptunian dwarf planet candidate, 2017 OF201, from the archival data of DECaLS. It is currently located at a distance of 90 au. When assuming a typical albedo of 0.13, its diameter is estimated to be 700 km, large enough to qualify as a dwarf planet.
@@
Its orbit is extremely wide and eccentric ($a=830$ au, $q=44.9$ au, $i=16.2^{\circ}$), influenced by both Neptune’s scattering and the Galactic tide at Gyr timescale, therefore representing an overlap between the scattering disk and the inner Oort cloud.
The probability for 2017 OF201 to be close enough and detectable is only 0.5% given its wide and eccentric orbit. Therefore, there should exist a population of hundreds of 2017 OF201-like objects in the outer solar system, totaling 1% of Earth’s mass, a significant fraction of the scattering disk.
The orbit of 2017 OF201 also poses a challenge to the Planet Nine (P9) hypothesis. Previous discoveries of extreme TNOs reveal an apparent clustering in longitude of perihelion near $\varpi\approx 60^{\circ}$, motivating the idea that an undetected planet shepherds their orbits. By contrast, 2017 OF201, with $\varpi=306^{\circ}$, lies well outside this clustering region (Figure 2) and appears dynamically inconsistent with confinement by P9. Our N-body simulations indicate that the inclusion of such a planet leads to ejection of 2017 OF201 in $\sim$60% of realizations, with a mean timescale of $\sim$0.3 Gyr (Figure 4). Further investigation will be necessary to assess whether this outcome robustly excludes the P9 hypothesis.
@@
Figure 5: Cutout images of all 22 detections from CFHT, DECam, and Magellan. Each image is 20 arcsec on a side. The detection time corresponds to the middle of each exposure. CFHT data are deeper and have a much better seeing than the DECam data. The detection spans a long time baseline from Aug 2011 to Jul 2025, enabling precise determination of its orbit.
Table 2: Archival and follow-up observations of 2017 OF201. The codes in parentheses are station codes of the observatory from the minor planet center. Note that the photometry reported here are measured in the native filter systems.
observation time R.A. rms Dec. rms magnitude band seeing expo. $\alpha$ $H(\alpha)$
UTC deg deg s deg
SDSS (645)
2004-09-21 06:02:01.31 20.4237305 0.08 +23.8643205 0.08 22.42 $\pm$ 0.12 $r$ 1.0 54 0.40 3.54
23.16 $\pm$ 0.42 $u$ 4.29
23.22 $\pm$ 0.16 $g$ 4.35
21.58 $\pm$ 0.09 $i$ 2.71
21.24 $\pm$ 0.25 $z$ 2.37
2009-09-26 10:49:07.74 22.9414929 0.10 +25.2131626 0.08 22.47 $\pm$ 0.12 $r$ 0.9 54 0.37 3.43
23.52 $\pm$ 0.84 $u$ 4.48
23.22 $\pm$ 0.17 $g$ 4.18
22.20 $\pm$ 0.14 $i$ 3.16
21.96 $\pm$ 0.41 $z$ 2.92
CFHT (T14)
2011-08-31 11:25:52.27 24.1672646 0.030 +25.7422073 0.030 22.659 $\pm$ 0.038 $r$ 0.9 320 0.59 3.543
2011-08-31 12:27:48.41 24.1669749 0.026 +25.7422438 0.025 22.604 $\pm$ 0.035 $r$ 0.9 320 0.59 3.488
2011-08-31 13:29:58.02 24.1666503 0.022 +25.7422514 0.022 22.634 $\pm$ 0.032 $r$ 0.8 320 0.59 3.518
2011-09-05 11:45:03.32 24.1300816 0.020 +25.7423370 0.018 22.582 $\pm$ 0.029 $r$ 0.7 320 0.55 3.467
2011-11-24 09:08:17.23 23.3120374 0.013 +25.4947651 0.013 22.588 $\pm$ 0.022 $r$ 0.5 320 0.38 3.471
2011-11-30 09:15:42.08 23.2618590 0.027 +25.4661271 0.026 22.610 $\pm$ 0.036 $r$ 0.9 320 0.44 3.490
2011-11-30 09:39:20.11 23.2617232 0.020 +25.4660430 0.020 22.610 $\pm$ 0.030 $r$ 0.7 320 0.44 3.490
2011-12-01 09:32:36.86 23.2539409 0.013 +25.4613167 0.013 22.632 $\pm$ 0.022 $r$ 0.5 320 0.45 3.512
2012-01-28 05:57:34.76 23.1285940 0.027 +25.2561656 0.024 22.656 $\pm$ 0.038 $r$ 0.8 384 0.68 3.508
DECam (W84)
2015-10-11 05:21:41.03 25.694931 0.15 +26.629005 0.18 23.32 $\pm$ 0.22 $g$ 1.5 125 0.25 4.08
2017-11-12 03:20:26.88 26.267165 0.19 +26.952323 0.22 24.32 $\pm$ 0.26 $g$ 1.3 200 0.24 5.01
2018-08-10 09:14:55.55 27.525852 0.07 +27.298418 0.08 23.49 $\pm$ 0.10 $g$ 1.3 168 0.66 4.14
2015-10-11 05:19:10.66 25.694897 0.10 +26.629028 0.10 22.54 $\pm$ 0.10 $r$ 1.3 119 0.25 3.23
2017-09-17 07:21:40.06 26.850032 0.06 +27.114610 0.07 22.71 $\pm$ 0.09 $r$ 1.2 110 0.46 3.40
2018-08-11 08:59:46.14 27.523248 0.05 +27.300693 0.05 22.62 $\pm$ 0.09 $r$ 0.9 40 0.66 3.27
2018-10-31 04:04:31.05 26.850075 0.11 +27.217787 0.12 22.65 $\pm$ 0.11 $r$ 1.8 118 0.18 3.31
2014-08-13 09:54:17.21 25.694468 0.05 +26.432920 0.05 21.79 $\pm$ 0.09 $z$ 0.8 146 0.67 2.57
2017-07-23 10:32:19.46 27.094235 0.11 +27.035966 0.09 21.77 $\pm$ 0.10 $z$ 1.8 250 0.68 2.44
2018-07-26 10:17:57.53 27.541774 0.11 +27.255624 0.10 22.02 $\pm$ 0.12 $z$ 1.4 229 0.67 2.67
Magellan (304)
2025-07-25 10:01:58.00 30.547527 0.10 +28.609305 0.10 22.69 $\pm$ 0.18 $i$ 1.0 600 0.64 3.11
2025-07-25 10:15:15.50 30.547569 0.05 +28.609342 0.05 23.10 $\pm$ 0.06 $r$ 1.1 600 0.64 3.52
2025-07-25 10:29:28.50 30.547500 0.12 +28.609326 0.12 23.79 $\pm$ 0.17 $g$ 1.2 700 0.64 4.21
@@ Appendix A Astrometry and Photometry
We perform astrometric and photometric measurements of the object using the CFHT/MegaCam, DECam, and Magellan imaging data. Cutouts of those images with 2017 OF201 centered are shown in Figure 5.
@@
For the detections in CFHT, we retrieve the single-epoch images from the Canada-France-Hawaii Telescope (CFHT)/MegaCam data archive [^16] https://www.cadc-ccda.hia-iha.nrc-cnrc.gc.ca/. These images are all in the $r$-band with 0.5–0.9 arcsec seeing. The instrument signatures (bias, dark, and flat) have been removed in these images, but they do not yet have astrometric solutions and photometric calibrations. We implement a two-step astrometric calibration process: first obtaining an initial solution using astrometry.net (Lang et al., 2010), followed by fine-tuning against Gaia DR3 reference stars (Gaia Collaboration et al., 2021) using SCAMP (Bertin, 2006). This procedure yields a final astrometric solution with an average positional uncertainty of $0.1\arcsec$. For photometric calibration, we first remove the sky background using SExtractor (Bertin and Arnouts, 1996) with a mesh size of $60\arcsec$. We then construct a point spread function (PSF) model using PSFEx (Bertin, 2011) from non-saturated point sources in each image, and perform PSF-fitting photometry using SExtractor. The photometric zeropoint is thus determined by comparing the instrumental magnitudes of non-saturated stars with their Pan-STARRS1 PSF magnitudes. Our object has an average $r$-band magnitude of $r=22.61$ mag in the CFHT data from 2011 to 2012.
@@
The DECam single-epoch images are downloaded from the Legacy Surveys [^17] https://www.legacysurvey.org/rawdata/, which are reduced using the DECam Community Pipeline [^18] https://noirlab.edu/science/data-services/data-reduction-software/csdc-mso-pipelines/pl206. These images are photometrically calibrated using Pan-STARRS1 (Chambers et al., 2016) and have a preliminary astrometric solution. To ensure consistency with our CFHT analysis, we refine the astrometric calibration using SCAMP against the Gaia DR3 catalog. Similar to the CFHT data above, for each DECam image, we subtract the sky background with a mesh size of $60\arcsec$, generate the PSF model, and perform PSF-fitting photometry. The resulting photometry have higher uncertainties since the DECam images are shallower and have poorer seeing conditions (0.8–2.0 arcsec) than the CFHT images.
We also include the SDSS detections on 2004 and 2009 for our orbital fitting. In fact, those detections are listed as two objects in SDSS’s catalog, with object ID 1237666274738438899 and 1237680073395798979, and misclassified as a star and a galaxy. This is because SDSS has only single-epoch observations on most part of its coverage and therefore cannot distinguish slowly-moving objects from static ones. As the SDSS catalog has a high accuracy of astrometric calibration with a systematic error below 0.03″(Pier et al., 2003), we directly use the astrometry ($r$-band) and photometry values in the catalog (Fukugita et al., 1996; Gunn et al., 1998; Lupton et al., 1999). Note that when considering photometry, one should use columns of psfMag instead of the default magnitude in the catalog due to the mis-classification of object type.
Appendix B Absolute magnitude
The phase angle $\alpha$ and reduced magnitude $H(\alpha)$ of each observations of 2017 OF201 are listed in table 2, using the best-fit orbit. The absolute magnitude $H$ is defined as the reduced magnitude $H(\alpha)=m-5\lg(r\Delta)$ when $\alpha=0$, where $\alpha$ is the phase angle, $m$ is the apparent magnitude, $r$ is the heliocentric distance, $\Delta$ is the geocentric distance. The phase curve $H(\alpha)$ reveals surface reflective properties of the object. Due to the limited range of phase angle of TNOs (0–2), phase curves are usually fit with a linear function $H(\alpha)=H+\beta\alpha$ (e.g. Rabinowitz et al., 2007).
For 2017 OF201, only the $r$-band has good enough photometry to fit simultaneously $H$ and $\beta$, which leads to $H_{r}=3.33\pm 0.05$ and $\beta_{r}=0.12\pm 0.11$ per degree. The slope $\beta$ is consistent with the average value of 0.15 measured from other TNOs (e.g. Sheppard and Jewitt, 2002; Rabinowitz et al., 2007), though the uncertainty is yet to be improved. The absolute magnitudes in other bands listed in table 2 are obtained by a least square fitting with the slope fixed to $\beta=0.12\pm 0.11$ per degree. Before fitting, all observed magnitudes from SDSS, CFHT, and Magellan (McLeod et al., 2015) have been converted into the DECam system [^19] https://des.ncsa.illinois.edu/releases/dr2/dr2-docs/dr2-transformations.